Point charges of 0.24 uC and 0.45 uC are placed 0.35 m apart. What is the magnitude of the electric field, in newtons per coulomb, halfway between them?
Answer #1

r = l/2 = 0.35 / 2 = 0.175 m

E = [kq2/r2] - [kq1/r2]

= (k / r2) (q2 - q1)

= (9 * 109 / 0.1752) * (0.45 * 10-6 - 0.24 * 10-6)

magnitude of the electric field​ = 6.17 * 104 N/C